# Scaling Generalist Data-Analytic Agents

### v1

@@ Abstract
Data-analytic agents are emerging as a key catalyst for automated scientific discovery and for the vision of Innovating AI. Current approaches, however, rely heavily on prompt engineering or multi-agent scaffolds over proprietary models, while open-source models still struggle with diverse-format, large-scale data files and long-horizon, multi-step reasoning that real-world analytics demands. This paper introduces DataMind, a scalable data synthesis and agent training recipe designed to construct generalist data-analytic agents. DataMind tackles three key challenges in building open-source data-analytic agents, including insufficient data resources, improper training strategy, and unstable code-based multi-turn rollout. Concretely, DataMind applies 1) a fine-grained task taxonomy and a recursive easy-to-hard task composition mechanism to increase the diversity and difficulty of synthesized queries; 2) a knowledge-augmented trajectory sampling strategy followed by model-based and rule-based filtering; 3) a dynamically adjustable training objective combining both SFT and RL losses; 4) a memory-frugal and stable code-based multi-turn rollout framework. Built on DataMind, we curate DataMind-12K, a high-quality trajectory set spanning diverse domains, task categories, and data file formats for data-analytic tasks. Trained on DataMind-12K, our DataMind-14B achieves state-of-the-art with an average score of 71.16% on multiple data analysis benchmarks, outperforming the strongest proprietary baselines DeepSeek-V3.1 and GPT-5. Our DataMind-7B also performs best among all open-source models with a score of 68.10%. We also list some empirical insights gained from our exploratory trials in the analysis experiments, aiming to provide actionable insights about agent training for the community. We will release DataMind-12K and DataMind-7B,14B for the community’s future research111  Code: https://github.com/zjunlp/DataMind..
@@
Figure 1: (a) Task Taxonomy. We categorize data analysis tasks into $18$ fine-grained categories to enhance the diversity of our synthesized queries. (b) Performance Comparison. Our DataMind-14B achieves the best compared with all proprietary models and open-source trained or untrained models.
@@ 2 Problem Definition
Our agent framework adheres to the prevailing ReAct (Yao et al., 2023) paradigm. Upon receiving a task, the agent is required to iterate multiple rounds of Thought–Action–Observation cycles and finally produce an answer. In the data analysis scenario, Thought denotes the agent’s reasoning and reflection process conditioned on the current context; Action refers to the agent’s invocation of code to process and compute over the data files or the generation of the final answer. The code may be written in 3Python or SQL, depending on the data file format; Observation consists of the execution feedback returned by the environment (i.e., ΘCode Interpreter).
@@ 3.1 File Collection and Query Synthesis
To generate specific queries, we devise an automated script to extract meta-information $d$ of each data file, such as table headers, column names, data types, and representative rows, and then feed these metadata into DeepSeek-V3 (DeepSeek-AI, 2024) to synthesize queries $q$. To ensure both diversity and fine-grainedness of the generated questions, we refer to and refine the taxonomy in Wu et al. (2025b) and classify the data analysis tasks into $18$ fine-grained categories (see Fig.1(a)). For each category, we carefully curate $4\sim 6$ exemplar queries that vary in complexity and domains and serve as few-shot demonstrations. Under the guidance of these type-specific contexts, every data file is used to generate a diverse set of queries that span the full spectrum of the proposed taxonomy. To further elevate query complexity, we adopt a recursive easy-to-hard composition scheme that chains multiple task types, i.e., the output of one task is fed as input to the next. By iterating $2\sim 5$ times, we progressively amplify the difficulty and create multi-hop analytic challenges that go well beyond the capability required by any single task type. The prompts for query synthesis can be found in Appx.G.2.
@@
To guarantee the quality of the synthesized trajectories, we introduce a knowledge-augmented trajectory sampling framework. Initially, for each question category, we manually craft a high-level workflow $k$ that encodes procedural knowledge and steers the model during trajectory synthesis. To further boost answer quality, we impose a self-consistency filter. We sample $\mathcal{N}$ independent trajectories per query and employ a judge model $\mathcal{M}$ powered by GPT-4o-mini (OpenAI, 2024b) to verify whether their final answers are consistent with reasoning rationales. Only trajectories that converge to the same answer are retained; among them, the judge model will also select the most concise and accurate one as our training instance $y$:
@@
where $c$ is the chain-of-thought process of the judge model to reach the binary conclusion $s$ of whether the sampled trajectories are consistent. We use DeepSeek-V3.1 (DeepSeek-AI, 2025) as our expert policy model $\pi_{\theta_{\textrm{expert}}}$. During implementation, we set $\mathcal{N}=3$. The prompt used for trajectory sampling and the judge model $\mathcal{M}$ can be found in Appx.G.3 and Appx.G.4, respectively. We extract the final answer from the trajectory as the final synthesized answer $a$ for the corresponding query $q$. However, this pipeline inherently biases us toward easier queries whose answers are more likely to coincide. To counteract this, we refine the high-level workflow knowledge $k$ into more granular, step-by-step instructions for categories that exhibit low inter-trajectory consistency. Moreover, for trajectories that fail the consistency check, we feed the judge model’s chain-of-thought back to the agent as external critique, prompting it to reflect and revise its reasoning path:
@@ 4 Scaling Data-Analytic Agent Training
Our reward mainly comprises three components: format reward $r_{\textrm{format}}$, answer reward $r_{\textrm{answer}}$, and length reward $r_{\textrm{length}}$. The agent is required to enclose its reasoning process within <think></think> tags, place any generated data-processing code between <code> and </code>, and wrap its final answer in <answer></answer>. The environment’s execution results will be placed between <interpreter> and </interpreter>. For the answer reward, as many answers are descriptive and thus resist rule-based verification, we adopt a model-as-judge powered by GPT-4o-mini (OpenAI, 2024b). We engineer a dedicated LLM evaluation prompt detailed in Appx. G.4. Both $r_{\textrm{format}}$ and $r_{\textrm{answer}}$ are binary with only 0 and $1$. To mitigate the risk of the agent hacking the answer reward by hallucinating excessive tokens, we further impose a length-based penalty to discourage overly verbose outputs. We define the length reward and the final reward as:
@@ 5 Experiments
Qwen-2.5
Coder-7B
ReAct 15.05 35.41 11.70 28.63 7.02 18.71 11.26 27.58
@@
Qwen-2.5
Coder-14B
ReAct 71.21 83.27 56.96 69.97 41.76 59.91 56.64 71.05
@@ 5.1 Experimental Settings
We evaluate our model on three datasets related to data analysis: DABench (Hu et al., 2024), TableBench (Wu et al., 2025b), and BIRD (Li et al., 2023b). Our evaluation protocol aligns with our answer reward method, where a judge model powered by GPT-4o-mini (OpenAI, 2024b) is used to evaluate the correctness of the final answer. We report both pass@1 and pass@3 scores for all the methods. Please refer to Appx.D for more details.
@@
We use LlamaFactory (Zheng et al., 2024) for SFT training and verl (Sheng et al., 2025) for RL training. For SFT, our learning rate is 1e51e-5 with a warmup ratio of $0.1$ and a cosine decay schedule. Our global batch size is set to $16$. For RL, we use a learning rate of 1e61e-6. The batch size is $16$ with a mini batch size of $2$. The rollout temperature is $0.7$, the top-p is $1.0$, and the group size $G$ is $4$. We schedule $\gamma$ via cosine decay, annealing from a peak of $0.9$ to a valley of $0.05$. At test time, we fix the temperature to $0.7$, top-p to $0.95$, and an inference batch size of 5 for all evaluations. The detailed hyperparameter information can be seen in Appx.F.
@@ 5.2 Main Results
As shown in Tab. 1, our 7B model, DataMind-7B, achieves the best among all open-source models with an average score of 68.10%. Our 14B model, DataMind-14B, attains an average score of 71.16% across all tasks, surpassing all proprietary models (including the latest GPT-5 and DeepSeek-v3.1) as well as all open-source alternatives. Moreover, our DataMind series models demonstrate robust mastery of diverse data formats and exhibit balanced performance across all datasets. By contrast, specialized models degrade sharply when confronted with unseen data. For example, OmniSQL-7B reaches 57.11% on BIRD, yet its performance on TableBench and DABench drops steeply. Note that to ensure a fair evaluation, we have converted all tables in these two benchmarks into .sqlite files. Nevertheless, SQL-oriented models still underperform. This observation indicates the breadth of query types and file formats covered by DataMind-12K. Furthermore, TableLLM and Table-R1 are limited to small-scale tables. When evaluated on DABench’s large-scale tables, they fail to generalize, and their accuracy deteriorates even further on BIRD’s multi-table analysis. These results highlight our model’s capacity to handle complex tabular data, which can be attributed to the difficulty distribution embedded in DataMind-12K. Moreover, all trained baselines are exposed to significantly larger training corpora than ours (20K instances for TableLLM and Table-R1, and 2.5M for OmniSQL and SQL-R1, versus only 12K for DataMind), yet we outperform them even on their adept benchmarks. This gain is attributable to the high-quality reasoning trajectories curated in DataMind-12K and our stable training strategy. Our model also maintains a high pass@3 score, indicating that it preserves strong generation diversity while ensuring reliability.
@@ 5.3 Analysis
Although SFT loss serves as an effective stabilizer for RL training, we find that its persistent dominance throughout training can conversely trigger collapse. As shown in Fig.5, fixing $\gamma$ at a high level also causes the answer reward to rise briefly, followed by a gradual decline. The underlying reason is that over-fitting to the SFT loss traps the policy in the rigid thinking patterns embedded in the expert trajectories, especially when
these trajectories are synthesized from the same model, thereby crippling exploration. To corroborate this, we track the entropy of the policy during training and observe a pronounced entropy collapse phenomenon. In contrast, our dynamic-$\gamma$ strategy can keep the policy entropy consistently at a relatively high level throughout training. Overall, we find the training process resembles raising a child. During early childhood, constant parental guidance (a large $\gamma$) is indispensable to keep the child from going astray. As the child grows up, excessive supervision stifles the child’s innate drive for self-directed exploration. At that stage, judiciously letting go (a small $\gamma$) enables the child to discover their true capabilities through the feedback from the surrounding world.
@@
Fig. 6 shows the impact of different degrees of cold start on the 7B model to subsequent RL training. For rapid empirical verification, we randomly sample $3,843$ training data from DataMind-12K (balanced on query types) and $240$ test data ($60$ for each of the three test datasets) for evaluation. As the number of cold start training epochs increases, the marginal gain achieved by RL over the cold start checkpoint (i.e., the slope of the dashed line) diminishes. This indicates that RL can narrow the performance gap between different base models (Liu et al., 2025). Nevertheless, although the gap is narrowed, post-RL performance remains positively correlated with the capability of the base model. This suggests that the bulk of knowledge is acquired during SFT, whereas RL primarily serves to unlock latent potential rather than explicitly push the model beyond its inherent capacity boundary (Yue et al., 2025a; Chu et al., 2025). Setting aside overfitting, the current trend suggests that a cross point may emerge in which a sufficiently strong cold start leaves no room for further improvement via RL. Whether such a point truly exists, and, if it does, what fundamental mechanisms (e.g., saturation of the policy space, diminishing exploratory signal, or intrinsic limitations of the reward model) render further RL ineffective, constitutes an important open question for future work.
@@ 7 Conclusion
We have submitted all our training and evaluation code in the Supplementary Material. Due to OpenReview’s file size limit, we only upload a $3,843$ subset of DataMind-12K. We will fully open DataMind-12K and our models DataMind-7B and 14B immediately after the double blind review process. The detailed training data synthesis and agent training methods can be found in §3 and §4. We have clearly reported the details of our evaluation datasets and metrics in §5.1 and Appx.D. The detailed information about the models and baselines we use, including the model versions of proprietary models and the reproduction details of baseline models, can be found in §5.1 and Appx.E. The code framework used and the training and inference hyperparameters are mentioned in §5.1 and Appx.F. All the prompts used in our paper are presented in Appx.G.

### v2

# arXiv:2509.25084v2[cs.CL] 27 Feb 2026

@@ Abstract
Data-analytic agents are emerging as a key catalyst for automated scientific discovery and for the vision of Innovating AI. Current approaches, however, rely heavily on prompt engineering or multi-agent scaffolds over proprietary models, while open-source models still struggle with diverse-format, large-scale data files and long-horizon, multi-step reasoning that real-world analytics demands. This paper introduces DataMind, a scalable data synthesis and agent training recipe designed to construct generalist data-analytic agents. DataMind tackles three key challenges in building open-source data-analytic agents, including insufficient data resources, improper training strategy, and unstable code-based multi-turn rollout. Concretely, DataMind applies 1) a fine-grained task taxonomy and a recursive easy-to-hard task composition mechanism to increase the diversity and difficulty of synthesized queries; 2) a knowledge-augmented trajectory sampling strategy followed by model-based and rule-based filtering; 3) a dynamically adjustable training objective combining both SFT and RL losses; 4) a memory-frugal and stable code-based multi-turn rollout framework. Built on DataMind, we curate DataMind-12K, a high-quality trajectory set spanning diverse domains, task categories, and data file formats for data-analytic tasks. Trained on DataMind-12K, our DataMind-14B achieves state-of-the-art with an average score of 71.16% on multiple data analysis benchmarks, outperforming the strongest proprietary baselines DeepSeek-V3.1 and GPT-5. Our DataMind-7B also performs best among all open-source models with a score of 68.10%. We also list some empirical insights gained from our exploratory trials in the analysis experiments, aiming to provide actionable insights about agent training for the community. We have released DataMind-12K and DataMind-7B,14B for the community111  Code: https://github.com/zjunlp/DataMind..
@@
Figure 1: (a) Task Taxonomy. We categorize data analysis tasks into $18$ fine-grained categories to enhance the diversity of our synthesized queries. (b) Performance Comparison. Our DataMind-14B achieves the best compared with all proprietary models and open-source trained or untrained models. * Note that DeepSeek-V3.1 is an open-sourced model. We categorize it as a proprietary model for representational convenience only.
@@ 2 Problem Definition
Our agent framework adheres to the prevailing ReAct (Yao et al., 2023) paradigm. Upon receiving a task, the agent is required to iterate multiple rounds of Thought–Action–Observation cycles and finally produce an answer. In the data analysis scenario, Thought denotes the agent’s reasoning and reflection process conditioned on the current context; Action refers to the agent’s invocation of code to process and compute over the data files or the generation of the final answer. The code may be written in Python or SQL, depending on the data file format; Observation consists of the execution feedback returned by the environment (i.e., Code Interpreter).
@@ 3.1 File Collection and Query Synthesis
To generate specific queries, we devise an automated script to extract meta-information $d$ of each data file, such as table headers, column names, data types, and representative rows, and then feed these metadata into DeepSeek-V3 (DeepSeek-AI, 2024) to synthesize queries $q$. To ensure both diversity and fine-grainedness of the generated questions, we refer to and refine the taxonomy in Wu et al. (2025b) and classify the data analysis tasks into $18$ fine-grained categories (see Fig.1(a)). For each category, we carefully curate $4\sim 6$ exemplar queries that vary in complexity and domains and serve as few-shot demonstrations. Under the guidance of these type-specific contexts, every data file is used to generate a diverse set of queries that span the full spectrum of the proposed taxonomy. To further elevate query complexity, we adopt a recursive easy-to-hard composition scheme that chains multiple task types, i.e., the output of one task is fed as input to the next. By iterating $2\sim 5$ times, we progressively amplify the difficulty and create multi-hop analytic challenges that go well beyond the capability required by any single task type. The prompts for query synthesis can be found in Appx.I.2.
@@
To guarantee the quality of the synthesized trajectories, we introduce a knowledge-augmented trajectory sampling framework. Initially, for each question category, we manually craft a high-level workflow $k$ that encodes procedural knowledge and steers the model during trajectory synthesis. To further boost answer quality, we impose a self-consistency filter. We sample $\mathcal{N}$ independent trajectories per query and employ a judge model $\mathcal{M}$ powered by GPT-4o-mini (OpenAI, 2024a) to verify whether their final answers are consistent with reasoning rationales. Only trajectories that converge to the same answer are retained; among them, the judge model will also select the most concise and accurate one as our training instance $y$:
@@
where $c$ is the chain-of-thought process of the judge model to reach the binary conclusion $s$ of whether the sampled trajectories are consistent. We use DeepSeek-V3.1 (DeepSeek-AI, 2025) as our expert policy model $\pi_{\theta_{\textrm{expert}}}$. During implementation, we set $\mathcal{N}=3$. The prompt used for trajectory sampling and the judge model $\mathcal{M}$ can be found in Appx.I.3 and Appx.I.4, respectively. We extract the final answer from the trajectory as the final synthesized answer $a$ for the corresponding query $q$. However, this pipeline inherently biases us toward easier queries whose answers are more likely to coincide. To counteract this, we refine the high-level workflow knowledge $k$ into more granular, step-by-step instructions for categories that exhibit low inter-trajectory consistency. Moreover, for trajectories that fail the consistency check, we feed the judge model’s chain-of-thought back to the agent as external critique, prompting it to reflect and revise its reasoning path:
@@ 4 Scaling Data-Analytic Agent Training
Our reward mainly comprises three components: format reward $r_{\textrm{format}}$, answer reward $r_{\textrm{answer}}$, and length reward $r_{\textrm{length}}$. The agent is required to enclose its reasoning process within <think></think> tags, place any generated data-processing code between <code> and </code>, and wrap its final answer in <answer></answer>. The environment’s execution results will be placed between <interpreter> and </interpreter>. For the answer reward, as many answers are descriptive and thus resist rule-based verification, we adopt a model-as-judge powered by GPT-4o-mini (OpenAI, 2024a). We engineer a dedicated LLM evaluation prompt detailed in Appx.I.4. Both $r_{\textrm{format}}$ and $r_{\textrm{answer}}$ are binary with only 0 and $1$. To mitigate the risk of the agent hacking the answer reward by hallucinating excessive tokens, we further impose a length-based penalty to discourage overly verbose outputs. We define the length reward and the final reward as:
@@ 5 Experiments
Qwen-2.5 Coder-7B ReAct 15.05 35.41 11.70 28.63 7.02 18.71 11.26 27.58
@@
Qwen-2.5 Coder-14B ReAct 71.21 83.27 56.96 69.97 41.76 59.91 56.64 71.05
@@ 5.1 Experimental Settings
We evaluate our model on three datasets related to data analysis: DABench (Hu et al., 2024), TableBench (Wu et al., 2025b), and BIRD (Li et al., 2023b). Our evaluation protocol aligns with our answer reward method, where a judge model powered by GPT-4o-mini (OpenAI, 2024a) is used to evaluate the correctness of the final answer. We report both pass@1 and pass@3 scores for all the methods. Please refer to Appx.D for more details.
@@
We use LlamaFactory (Zheng et al., 2024) for SFT training and verl (Sheng et al., 2025) for RL training. For SFT, our learning rate is 1e51e-5 with a warmup ratio of $0.1$ and a cosine decay schedule. Our global batch size is set to $16$. For RL, we use a learning rate of 1e61e-6. The batch size is $32$ with a mini batch size of $4$ and the number of epochs is 1. The rollout temperature is $0.7$, the top-p is $1.0$, and the group size $G$ is $4$. We schedule $\gamma$ via cosine decay, annealing from a peak of $0.9$ to a valley of $0.05$. At test time, we fix the temperature to $0.7$, top-p to $0.95$, and an inference batch size of 5 for all evaluations. For all the processes, the maximum number of interaction rounds $\mathcal{T}$ is set to 10. Each of the training experiments can be run on a machine with 8 80G A100 GPUs within 2 days. The detailed hyperparameter information can be seen in Appx.F.
Figure 3: Ablation on Data.
@@ 5.2 Main Results
As shown in Tab.1, our 7B model, DataMind-7B, achieves the best among all open-source models with an average score of 68.10%. Our 14B model, DataMind-14B, attains an average score of 71.16% across all tasks, surpassing all proprietary models (including the latest GPT-5 and DeepSeek-v3.1) as well as all open-source alternatives. Moreover, our DataMind series models demonstrate robust mastery of diverse data formats and exhibit balanced performance across all datasets. By contrast, specialized models degrade sharply when confronted with unseen data. For example, OmniSQL-7B reaches 57.11% on BIRD, yet its performance on TableBench and DABench drops steeply. Note that to ensure a fair evaluation, we have converted all tables in these two benchmarks into .sqlite files. Nevertheless, SQL-oriented models still underperform. This observation indicates the breadth of query types and file formats covered by DataMind-12K. Furthermore, TableLLM and Table-R1 are limited to small-scale tables. When evaluated on DABench’s large-scale tables, they fail to generalize, and their accuracy deteriorates even further on BIRD’s multi-table analysis. These results highlight our model’s capacity to handle complex tabular data, which can be attributed to the difficulty distribution embedded in DataMind-12K. Moreover, all trained baselines are exposed to significantly larger training corpora than ours (20K instances for TableLLM and Table-R1, and 2.5M for OmniSQL and SQL-R1, versus only 12K for DataMind), yet we outperform them even on their adept benchmarks. This gain is attributable to the high-quality reasoning trajectories curated in DataMind-12K and our stable training strategy. Our model also maintains a high pass@3 score, preserving strong generation diversity while ensuring reliability.
@@
5.3 Ablation Studies
Ablation on Data Volume.
We conduct an ablation study on the impact of training data scale by randomly sampling 2K, 4K, and 8K instances from DataMind-12K to train the agent. The average performance of the 7B model on all the benchmarks is reported in Fig.3. It shows that as the training data volume increases, the model’s performance exhibits a clear scaling law, demonstrating the scalability of the DataMind data-synthesis pipeline. Another intriguing observation is that the gap between pass@1 and pass@3 gradually narrows as the data size grows. This pattern seems to align with the conclusion in Yue et al. (2025a): RL training tends to increase the likelihood of generating correct samples, yet contributes little to expanding the model’s coverage of solvable problems.
Table 2: Ablation Results of Different Agent Training Strategy.
Method Avg.
pass@1 pass@3
SFT 62.54 73.74
zero-RL 58.03 71.72
SFT-then-RL 63.42 75.46
SFT-and-RL 68.10 79.07
Ablation on Training Strategy.
In Tab.2, we demonstrate the superiority of our hybrid SFT+RL training objective by comparing with pure SFT, zero-RL, and SFT-then-RL on the 7B model. We need to say that the majority of our performance gains stem from our high-quality training data, as SFT alone raises the score from 11.26% (ReAct) to 62.54%, substantially closing the gap with other strong models. However, zero-RL performs markedly worse than SFT, while SFT-then-RL yields only marginal gains over pure SFT. On the contrary, our hybrid training strategy can further boost the model performance to 68.10% of pass@1 and 79.07% of pass@3. What these numbers do not reveal is the training stability conferred by our hybrid objective, as both zero-RL and SFT-then-RL can easily fall into collapse (detailed analysis about training stability can be found in §5.4), and we need to run many trials and select a relatively good checkpoint that performs best on the validation set.
Figure 5: Ablation on Judge Model. Solid markers denote results on DABench; hollow markers correspond to TableBench.
Ablation on Evaluation Method.
Because the same judge model is used during both training and evaluation, to rule out the possibility that our model wins by hacking the judge, we re-score representative baselines on DABench and TableBench with Qwen-2.5-72B serving as an external judge, and plot the pairwise mapping between the two judges in Fig.5. As shown, the rankings produced by the two judges are virtually identical (each model–dataset pair contributes two points, corresponding to pass@1 and pass@3). Quantitatively, the Pearson correlation between the two sets of scores is 0.96, indicating an almost perfect linear relationship. Furthermore, we isolate the descriptive-generation tasks in TableBench and evaluate the outputs of DataMind and the baselines with the dataset’s original objective metric, Rouge-L, in Appx.G.4. The resulting rankings again align closely with those produced by our model-as-judge. Nevertheless, Rouge-L clearly fails to capture true model performance. It over-emphasizes surface lexical and sentence overlap with the gold label rather than answer correctness. Even the powerful DeepSeek-V3.1 attains a very low Rouge-L score, underscoring the rationality and fairness of employing a judge model for evaluation.
@@ 5.4 Analysis
Although SFT loss serves as an effective stabilizer for RL training, we find that its persistent dominance throughout training can conversely trigger collapse. As shown in Fig.7, fixing $\gamma$ at a high level also causes the answer reward to rise briefly, followed by a gradual decline. The underlying reason is that over-fitting to the SFT loss traps the policy in the rigid thinking patterns embedded in the expert trajectories, especially when these trajectories are synthesized from the same model, thereby crippling exploration. To corroborate this, we track the entropy of the policy during training and observe a pronounced entropy collapse phenomenon. In contrast, our dynamic-$\gamma$ strategy can keep the policy entropy consistently at a relatively high level throughout training. Overall, we find the training process resembles raising a child. During early childhood, constant parental guidance (a large $\gamma$) is indispensable to keep the child from going astray. As the child grows up, excessive supervision stifles the child’s innate drive for self-directed exploration. At that stage, judiciously letting go (a small $\gamma$) enables the child to discover their true capabilities through the feedback from the surrounding world.
@@
Fig.8 shows the impact of different degrees of cold start on subsequent RL training. We randomly sample $3,843$ training data from DataMind-12K (balanced on query types) and $240$ test data ($60$ for each of the three test datasets) for evaluation. As the number of cold start training epochs increases, the marginal gain achieved by RL over the cold start checkpoint (i.e., the slope of the dashed line) diminishes. This indicates that RL can narrow the performance gap between different base models (Liu et al., 2025). This phenomenon also aligns well between the 7B and 14B models with the same training epochs. Nevertheless, although the gap is narrowed, post-RL performance remains positively correlated with the capability of the base model. This suggests that the bulk of knowledge is acquired during SFT, whereas RL primarily serves to unlock latent potential rather than explicitly push the model beyond its inherent capacity boundary (Yue et al., 2025a; Chu et al., 2025).
@@ 7 Conclusion
We have submitted all our training and evaluation code in the Supplementary Material. Due to OpenReview’s file size limit, we only upload a $3,843$ subset of DataMind-12K. We will fully open DataMind-12K and our models DataMind-7B and 14B immediately after the double blind review process. The detailed training data synthesis and agent training methods can be found in §3 and §4. We have clearly reported the details of our evaluation datasets and metrics in §5.1 and Appx.D. The detailed information about the models and baselines we use, including the model versions of proprietary models and the reproduction details of baseline models, can be found in §5.1 and Appx.E. The code framework used and the training and inference hyperparameters are mentioned in §5.1 and Appx.F. All the prompts used in our paper are presented in Appx.I.
Acknowledgments
We would like to express our sincere gratitude to the anonymous reviewers for their thoughtful and constructive feedback. This work was supported by the National Natural Science Foundation of China (No. 62576307, No. NSFCU23B2055, No. NSFCU19B2027), the Fundamental Research Funds for the Central Universities (226-2023-00138), the 2025 Zhejiang Provincial Center for Disease Control and Prevention Science and Technology Talent Incubation Project (No. 2025-A-04), the 2025 Zhejiang Health Informatics Association Scientific Research Program (Key Project, No. 2025XHZN-Z01), titled “Research on Monitoring and Early Warning Methods of AI Large Model and Infectious Disease Epidemic Data Fusion”, undertaken by the Zhejiang Provincial Center for Disease Control and Prevention, the Yongjiang Talent Introduction Programme (2021A-156-G), and the Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.